The present age of father exceeds 3 times the present age of his son by 2 years. six years ago, the age of the father exceeded 4 times the age of the son by 4 years. Find the present age of the son and father.
Answer by ankor@dixie-net.com(22740)

The present age of father exceeds 3 times the present age of his son by 2 years.
six years ago, the age of the father exceeded 4 times the age of the son by 4 years.
Find the present age of the son and father.
:
let f = father's present age
let s = son's present age
:
Write an equation for each statement:
:
"The present age of father exceeds 3 times the present age of his son by 2 years."
f = 3s + 2
:
" six years ago, the age of the father exceeded 4 times the age of the son by 4 years."
f - 6 = 4(s-6) + 4
f - 6 = 4s - 24 + 4
f - 6 = 4s - 20
f = 4s - 20 + 6
f = 4s - 14
replace f with (3s+2) from the 1st statement:
3s + 2 = 4s - 14
2 + 14 = 4s - 3s
16 = s
son is 16 yrs old
then
f = 3(16) + 2
f = 48 + 2
f = 50 yrs is the father's age
:
:
Check solution in the 2nd statement:
" six years ago, the age of the father exceeded 4 times the age of the son by 4 years."
50 - 6 = 4(16 - 6) + 4
44 = 4(10) + 4; confirms our solutions